### v1

# Charged excitations made neutral: N-centered ensemble density functional theory of Fukui functions

Abstract
We introduce an in-principle exact working equation to compute Fukui functions within $N$-centered ensemble DFT. It avoids the kernel derivative discontinuity problem of DFT for fractional number of electrons, whose contribution is recovered through weight-derivatives of the ensemble density functional potential. We explore practical strategies to compute its contribution by recycling ground-state density-functional approximations dressed with a weight-dependent scaling function. We also show interpolating between known limits of the ensemble functional and learning from uniform density profiles is a very effective strategy.
Introduction – No matter how much we succeed in making accurate computer simulations of chemical reactions routine, there will always be a need for chemical descriptors of reactivity. While they cannot systematically replace simulation nor experiment, their appeal is to bring ”primitive patterns of understanding” [^1] Attributed to C.A. Coulson by R. McWeeny in the preface of McWeeny R (1979) Coulson’s valence. Oxford University Press, Oxford reducing the over-abundant information generated in Ab Initio resolution of so many processes to key principles we can grasp. In a time where accurate molecular simulations are not so routinely feasible, their guidance is as valuable at the prediction stage than for ensuing rationalization.
As chemistry and condensed-matter physics is first and foremost dictated by the response of electrons to perturbation (from other electrons and nuclei, light, applied electric field and many more), they are the focus of properties descriptors. The advent of Density Functional Theory (DFT), an in-principle exact framework in which every property of the system follows from the electronic density, brought a solid foundation from which sharp definitions of insightful concepts – such as electronegativityParr (1980) and hardnessParr and Pearson (1983) – emerged. Further analysis of increasingly local responses of the electronic energy to changes in the external potential and number of electrons bloomed into the field of Conceptual Density Functional TheoryGeerlings et al. (2003); Geerlings and De Proft (2008) (CDFT). In turn, their computation from corresponding functional derivatives of the (to-be-approximated) universal functional proved instrumental in understanding what information practical Density Functional Approximations (DFA) should incorporate: By focusing efforts on reproducing more relevant properties, DFA design can (hopefully) find shortcuts on the meandering path towards the exact functionalGould (2017). This is a contribution of the present work.
We will focus more specifically on the Fukui function, introduced by Parr and YangParr and Yang (1984) as a tool to understand the relative reactivity of different sites in a moleculeAyers and Levy (2000). It is defined as the derivative of the electronic density with respect to the electron number, a somewhat puzzling concept for isolated moleculesNguyen-Dang et al. (1982) which must have an integer number of electrons, but very natural for atoms or fragments within a bigger molecular arrangement resulting in partial charges on each subunitBader (1990). Based on the reasoning that an approaching reagent will prefer the site most favorable to addition/removal of an electron for a nucleophilic/electrophilic attack, the corresponding Fukui function will reach a higher value the more reactive a site isAyers and Levy (2000). It offers a (sometimes criticalBartolotti and Ayers (2005)) improvement over frontier molecular orbital theoryFukui et al. (1952, 1954); Fukui (1982) by inclusion of orbital relaxation effects.
Interpretation and computation of derivatives with respect to the number of electrons analytically, i.e. connecting them to the $N$-electron Kohn Sham system of DFT and the Hartree-exchange-correlation (Hxc) density functional, is much more insightful (and more accurate) than their finite-differences evaluation from separate calculations on the $N$ and NN-1 electron systemsYang et al. (2012). To that end, extension of DFT to fractional electron numbers by Perdew, Parr, Levy and Balduz (PPLB) showed the generalized energy functional possesses derivative discontinuities when crossing an integer number of electronsPerdew et al. (1982), a behavior which in turn affects properties obtained through differentiating itPerdew et al. (1982); Mori-Sánchez et al. (2008); Cohen et al. (2008a, b); Stein et al. (2010); Zheng et al. (2011); Cohen et al. (2011); Perdew et al. (2017). Accurate predictions depend on inclusion of discontinuous shifts of the Hxc fields at integer electron number thresholds, which practical (local and semi-local) DFAs lack. The most well-known case is the missed contribution of Hxc potential derivative discontinuities resulting in systematic underestimation of fundamental gaps in solidsPerdew et al. (2017), also relevant for molecular dissociation and charge transferHodgson et al. (2017); Kraisler et al. (2021). It is less widely appreciated that similar discontinuities also appear in the Hxc kernel, responsible for aforementioned orbital relaxation effects, making computation of the Fukui function within PPLB challengingHellgren and Gross (2012). Constraints on KS orbital densities arising from piecewise-linearity requirement of the physical density were recently investigatedKraisler (2025), aiming to identify and correct density-driven errors in DFT calculations. The HOMO dependence on its fractional occupation proved to be intricate, invalidating the frozen approximation and even linear behavior in simple atomic systems. In some cases, it was found to not even be two-point-Taylor-expandable, with preliminary analysis suggesting the divergence of Taylor coefficients stems from approximate functionals being used.
The recently introduced $N$-centered extension of ensemble density functional theorySenjean and Fromager (2018); Hodgson et al. (2021); Cernatic et al. (2022) (Nc EDFT) is an alternative exact approach to charged electronic excitations that shed a new light on the discontinuity problem. While PPLB considers an ensemble of $N$ and N1N-1 electons pure states with relative weights constraining the obtained ensemble density, the Nc ensemble density always integrates to the central number of electrons $N$ by construction, no matter the weights (vide infra). The ensemble density and weights become independent variables of the theory, which transform the elusive energy derivative discontinuity contributions into weight-derivatives of the Hxc ensemble functional. This offers a promising way to ”recycle” local KS DFA lacking derivative discontinuities. It also provides a considerable gain in scope as it can describe any charged or neutral excitation within a unified pictureCernatic et al. (2024), and can be adapted to treat open systemsSenjean and Fromager (2020). We can benefit from progress on DFA development for EDFT in recent yearsGould and Pittalis (2019); Fromager (2020); Gould et al. (2020, 2023); Gould and Pittalis (2024); Gould et al. (2025a, b). Moreover, the introduction by the Authors of in principle exact expressions for neutrally-excited state densitiesFromager (2025) and responsesDupuy and Fromager (2025) within an ensemble makes computation of properties beyond excitation energies a timely pursuit.
In this Letter, we derive an exact $N$-centered EDFT of Fukui functions and discuss the fundamental and practical implications of our findings in the description of second (and possibly higher)-order density-functional Hxc derivatives. We then investigate two strategies to build a practical EDFAs and study their performance on the Hubbard dimer model systemCarrascal et al. (2015) as proof of principle. First, we will show how dressing ground-state KS DFAs with a weight-dependent scaling function allows to capture the weight-derivative term. Then, we will demonstrate the performance of an EDFA built by interpolating between known low and high correlation limits of the Hxc kernel, with additional constraint from the ground-state uniform electron gas limit.
[…]
$n^{\bm{\xi}}({\bf r})=\left(1-\sum_{\lambda>0}\frac{N_{\lambda}}{N}\xi_{\lambda}\right)n_{\Psi_{0}}({\bf r})+\sum_{\lambda>0}\xi_{\lambda}n_{\Psi_{\lambda}}({\bf r}),$ (2)
where $n_{\Psi}(\mathbf{r}):=\langle\Psi|\hat{n}(\mathbf{r})|\Psi\rangle$, integrates to the reference (so-called central) electron number $N$ of the reference $\nu=0$ ground state. It thus generalizes the Theophilou-Gross-Oliveira-Kohn (TGOK) ensemble framework for neutral excitations Theophilou (1979); Gross et al. (1988a, b); Oliveira et al. (1988) to also describe charged excitations, and has already been explored in some detailSenjean and Fromager (2018); Deur and Fromager (2019); Hodgson et al. (2021); Cernatic et al. (2022); Cernatic and Fromager (2024); Cernatic et al. (2024), in particular considering consequences of an ensemble density $n$ and $\bm{\xi}$ being independent variables of the theory, unlike in regular (PPLB) DFT Cernatic et al. (2022). In the conventional KS formulation of Nc EDFT, the ensemble density is mapped onto a non-interacting density matrix operator $\hat{\Gamma}_{s}^{\bm{\xi}}=\left(1-\sum_{\lambda>0}\frac{N_{\lambda}}{N}\xi_{\lambda}\right)|{\Phi^{\bm{\xi}}_{0}}\rangle\langle{\Phi^{\bm{\xi}}_{0}}|+\sum_{\lambda>0}\xi_{\lambda}|{\Phi^{\bm{\xi}}_{\lambda}}\rangle\langle{\Phi^{\bm{\xi}}_{\lambda}}|$, i.e.,
[…]
$n_{\Psi_{\nu}}=\frac{N_{\nu}}{N}n^{\bm{\xi}}+\sum_{\lambda>0}(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda})\frac{\partial n^{\bm{\xi}}}{\partial\xi_{\lambda}}$.
According to the KS mapping of Eq. (3), it leads to the more explicit expression (where we dropped the $\mathbf{r}$ dependence, for the sake of compactness),
$n_{\Psi_{\nu}}=n_{\Phi^{\bm{\xi}}_{\nu}}+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\chi^{\bm{\xi}}_{s}\star\dfrac{\partial}{\partial\xi_{\lambda}}\left(v^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]\right),$ (4)
where we use the shorthand notation $\chi\star f:\mathbf{r}\mapsto\int d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime})f(\mathbf{r}^{\prime})$ and $\chi^{\bm{\xi}}_{s}$ is the non-interacting KS analog of the Nc ensemble density-density linear response function $\chi^{\bm{\xi}}:\,(\mathbf{r},\mathbf{r}^{\prime})\mapsto\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta n^{\bm{\xi}}(\mathbf{r})/\delta v(\mathbf{r}^{\prime})$. Note that the two response functions are related through the following Dyson equation,
$\big(\chi^{\bm{\xi}}\big)^{-1}=\big(\chi^{\bm{\xi}}_{s}\big)^{-1}-f^{\bm{\xi}}_{\rm Hxc},$ (5)
where $f^{\bm{\xi}}_{\rm Hxc}=f^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]$ is the Nc ensemble density-functional Hxc kernel $f^{\bm{\xi}}_{\rm Hxc}[n](\mathbf{r},\mathbf{r}^{\prime})\equiv\delta v^{\bm{\xi}}_{\rm Hxc}[n](\mathbf{r})/\delta n(\mathbf{r}^{\prime})$ evaluated at the exact ensemble density. If we further expand and rewrite Eq. (4) as follows,
$\begin{split}&\big(\chi^{\bm{\xi}}_{s}\big)^{-1}\star\left(n_{\Psi_{\nu}}-n_{\Phi^{\bm{\xi}}_{\nu}}\right)=f^{\bm{\xi}}_{\rm Hxc}\star\left(n_{\Psi_{\nu}}-\frac{N_{\nu}}{N}n^{\bm{\xi}}\right)\\ &+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}},\end{split}$ (6)
or, equivalently,
$\begin{split}\big(\chi^{\bm{\xi}}\big)^{-1}\star n_{\Psi_{\nu}}&=\big(\chi^{\bm{\xi}}_{s}\big)^{-1}n_{\Phi^{\bm{\xi}}_{\nu}}-\frac{N_{\nu}}{N}f^{\bm{\xi}}_{\rm Hxc}\star n^{\bm{\xi}}\\ &\quad+\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}},\end{split}$ (7)
we finally obtain, by noticing that $\chi^{\bm{\xi}}\big(\chi^{\bm{\xi}}_{s}\big)^{-1}=1+\chi^{\bm{\xi}}\star f^{\bm{\xi}}_{\rm Hxc}$, the following exact and compact relation between the true and KS individual-state densities,
$\begin{split}n_{\Psi_{\nu}}\underset{\nu\geq 0}{=}&\big(1+\chi^{\bm{\xi}}\star f_{\rm Hxc}^{\bm{\xi}}\big)\star n_{\Phi^{\bm{\xi}}_{\nu}}-\frac{N_{\nu}}{N}\chi^{\bm{\xi}}\star f_{\rm Hxc}^{\bm{\xi}}\star n^{\bm{\xi}}\\ &+\chi^{\bm{\xi}}\star\sum_{\lambda>0}\Big(\delta_{\lambda\nu}-\frac{N_{\nu}}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}.\end{split}$ (8)
Note that Eq. (8) generalizes Eq. (104) of Ref. 41 to charged excited states. As it applies universally to any $N_{\nu}$-electron ground or excited state ($N_{\nu}=N,N\pm 1,N\pm 2,\ldots$), it can be (straightforwardly) applied to the calculation of physical (interacting) Fukui functions $f_{\kappa}={\theta}_{\kappa}(n_{\Psi_{\kappa}}-n_{\Psi_{0}})$, for which $N_{\kappa}\neq N$ and ${\theta}_{\kappa}={\rm sgn}(N_{\kappa}-N)$, thus leading to the following exact relation between $f_{\kappa}$ and its non-interacting KS analog $f^{\bm{\xi}}_{s,\kappa}={\theta}_{\kappa}(n_{\Phi^{\bm{\xi}}_{\kappa}}-n_{\Phi^{\bm{\xi}}_{0}})$,
[…]
Eq. (9) is the first key result of this Letter.
Interestingly, in regular (PPLB) DFT, this relation reduces to the first term on the right-hand side (the second and third terms simply do not exist in PPLB), which is then dependent on the aforementioned discontinuous Hxc kernel for accurate predictions Hellgren and Gross (2012). It appears that, within the Nc formalism, the impact of discontinuities on the calculation of Fukui functions is described explicitly through ensemble weight derivatives, as readily seen from the last term on the right-hand side of Eq. (9). Thus, we generalize a feature that was already known for the evaluation of ionization potentials and electron affinities Cernatic et al. (2024). This equivalence can be made even more transparent by using the invariance of Eq. (9) under symmetric shifts
$f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})\rightarrow f_{\rm Hxc}^{\bm{\xi}}({\bf r},{\bf r}^{\prime})+g({\bf r})+g({\bf r}^{\prime}),$ (10)
since $|\int d\mathbf{r}\,f^{\bm{\xi}}_{s,\kappa}(\mathbf{r})=\absolutevalue{N_{\kappa}-N}$ and $\int d\mathbf{r}\,\chi^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})=\delta\left(\int d\mathbf{r}\,n^{\bm{\xi}}(\mathbf{r})\right)/\delta v(\mathbf{r}^{\prime})=\delta N/\delta v(\mathbf{r}^{\prime})$=0. This fundamental property is specific to Nc EDFT. Indeed, in PPLB, physically-relevant variations of the density lead to a change of particle number, thus breaking the gauge-invariance Hellgren and Gross (2012). Note that, within Nc EDFT, we can always make a gauge choice for a specific charged excitation $0\rightarrow\kappa]$ such that the last two terms on the right-hand side of Eq. (9) cancel out, i.e.,
$\begin{split}&\chi^{\bm{\xi}}\star\bigg[\frac{|N_{\kappa}-N|}{N}f_{\rm Hxc}^{\bm{\xi}[0\rightarrow\kappa]}\star n^{\bm{\xi}}\\ &-{\theta}_{\kappa}\sum_{\lambda>0}\Big(\delta_{\lambda\kappa}-\frac{N_{\kappa}-N}{N}\xi_{\lambda}\Big)\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}\bigg]=0,\end{split}$ (11)
so that Eq. (9) becomes formally identical to that of PPLB, with no approximations made. Therefore, each excitation process requires applying a specific ($\kappa$-dependent) weight-dependent symmetric shift $g^{\bm{\xi}[0\rightarrow\kappa]}({\bf r})+g^{\bm{\xi}[0\rightarrow\kappa]}({\bf r}^{\prime})=f_{\rm Hxc}^{\bm{\xi}[0\rightarrow\kappa]}(\mathbf{r},\mathbf{r}^{\prime})-f_{\rm Hxc}^{\bm{\xi}}(\mathbf{r},\mathbf{r}^{\prime})$ to the ensemble Hxc kernel to ensure a systematic exactification of the PPLB formula, which then holds not only for ground but also for excited $N_{\kappa}$-electron states. According to Eq. (11),
$\begin{split}g^{\bm{\xi}[0\rightarrow\kappa]}&=\dfrac{1}{N_{\kappa}-N}\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\kappa}}\Big|_{n=n^{\bm{\xi}}}\\ &\quad-\dfrac{1}{N}\left(\sum_{\lambda>0}\xi_{\lambda}\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|_{n=n^{\bm{\xi}}}\right)-f_{\rm Hxc}^{\bm{\xi}}\star\dfrac{n^{\bm{\xi}}}{N}.\end{split}$ (12)
Eq. (12), which is our second and central formal result, establishes the equivalence between the discontinuities in the Hxc kernel and the weight derivatives of the ensemble Hxc density-functional potential (see the first term on the right-hand side). The clarity of the Nc EDFT picture is made manifest in the fact that no assumption about the asymptotic behavior of the density needs to be invoked Hellgren and Gross (2012) to rationalize and evaluate the shift. This is relevant for analysis of periodic systems, where Fukui functions can rationalize surface reactivity Barrera et al. (2025).
Practical strategies for EDFA design – By exploring the performance of EDFAs for computing Fukui functions through Eq. (9), we will discuss what a good approximation to the ensemble Hxc functional is, gather insights on the interplay of density-dependence and weight-dependence, and particularly, how Nc EDFT could possibly recycle semi-local standard DFAs whose lack of discontinuous behavior impeded PPLB. We do so by applying the Nc formalism to the Hubbard Dimer as it allows to study all regimes of electronic correlation (controlled by the parameter $U/t$) while the Hxc functional and its derivatives can be computed exactlyCarrascal et al. (2015). Note the strategies are general, and applicable to Ab Initio systems. We will focus on charged excitations between different electronic ground states, considering an ensemble of the $N=2$-electron state and the ”neighboring” $N\pm 1$-electron states with weights $\bm{\xi}=(\xi_{-},\xi_{+})$. In all results shown, we evaluate density functionals at the exact ensemble density. The ensemble response function $\chi^{\bm{\xi}}$ will be computed from Eq. (5) using the exact KS response.
The opportunity to employ regular (ground-state) DFAs arises from noticing that Eq.(9) holds even in the zero weights limit, where the ensemble Hxc kernel, response and density become those of regular $N$-electron ground-state DFT:
$\begin{split}f_{\pm}&=\big(1+\chi\star f_{\rm Hxc}\big)\star f_{s,\pm}-\;\chi\star f_{\rm Hxc}\star\frac{n_{0}}{N}\\ &\quad\pm\;\chi\star\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n_{0}}^{\bm{\xi}=0},\end{split}$ (13)
$f_{s,+}(\mathbf{r})$ is the lowest unoccupied KS molecular orbital (LUMO) in the $N$-electron ground state and $f_{s,-}(\mathbf{r})$ the highest occupied one (HOMO). Note $n_{0}=n^{\bm{\xi}=0}=n_{\Psi_{0}}$.
This is an astounding simplification: the only piece needed to compute Fukui functions aside from a ground-state functional without any approximation is the weight-derivatives the ensemble Hxc potential evaluated for null weights. In practice, one does not need to build an ensemble DFA covering the whole range of valid weights to obtain accurate results, but merely to model the weight dependence of the potential in the immediate vicinity of the familiar $N$-electron ground-state limit.
To that aim, it is appealing to ”dress” a regular KS-DFA with a weight-dependent scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s(\bm{\xi})\;v^{\bm{\xi}=0}_{\rm Hxc}[n]=s(\bm{\xi})\;v_{\rm Hxc}[n]$ (14)
where $s(\bm{\xi}=0)=1$ by construction. The weight-derivatives would then be approximated as
$\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\lambda}}\Big|^{\bm{\xi}=0}_{n=n_{0}}\approx\frac{\partial s(\bm{\xi})}{\partial\xi_{\lambda}}\Big|_{\bm{\xi}=0}v_{\rm Hxc}[n_{0}]$ (15)
This takes full advantage of the 0-weight limit mindset, $s(\bm{\xi})$ only accounting for the ensemble behavior in the vicinity of $\bm{\xi}=0$, and benefits from all the known exact constraints built in the ground-state functional provided the scaling function is chosen in a way that complies with them. In particular, this choice complies with exact uniform coordinate scaling constraint valid both for regular KS DFA and for ensembles Nagy (1995); Scott et al. (2024). Known limits of the ensemble Hxc functional can be used to build $s(\bm{\xi})$, as we will explore in the following. Throughout this work, we will use the exact 2-electrons Hubbard ground-state Hxc functional in pair with scaling functions.
Low correlation regime – Orbital-based approximations were used extensively in EDFT to bypass development of explicit (weight-dependent) density functionals, the simplest being an Hartree-Exchange approximation to the ensemble energyYang et al. (2014); Pribram-Jones et al. (2014); Yang et al. (2017); Gould et al. (2018); Sagredo and Burke (2018); Deur et al. (2018); Senjean and Fromager (2018); Deur and Fromager (2019); Senjean and Fromager (2020). When the correlation is not too strong, one can expect such an approach to do well, which we put to the test by considering the exact exchange ensemble functional (EEXX) of the Hubbard Dimer. We recall it is obtained as the first order term in the Taylor expansion in $U$ of the universal functional,
$E^{\bm{\xi}}_{\rm Hx}[n]=U\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}$ (16)
and yields all quantities in Eq. (9) by differentiation of $E^{\bm{\xi}}_{\rm Hx}[n]$. All results are computed in the 0-weight limit.
Figure 1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset). Full and dashed lines are $f_{+}$ and $f_{-}$ calculations respectively. Exact results (yellow lines) are compared to EEXX (blue lines) and to the neglect of weight-derivative contributions (grey lines).
In the weak correlation regime $U/t=0.75$, the EEXX functional closely reproduces exact results as seen on Fig. 1, with the effect of neglecting Hxc weight derivatives shown in grey. The improvement is clear, making this first application of EEXX to compute densities promising.
These encouraging results suggest using the scaling function
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hxc}[n]\,,\qquad s_{\rm Hx}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm Hx}[n^{\bm{\xi}}]}{v_{\rm Hx}[n^{\bm{\xi}}]}$ (17)
Note the ensemble density is fed to the ground-state potential. It is important to define the scaling in this way to ensure evaluating the weight-derivative of the scaling function (in the 0-weight limit) does not require knowing the weight-derivative of the ensemble density. While using this scaling function misses correlation at the ensemble level, note some measure of it is still brought into the approximate potential through the ground-state Hxc potential (see Eq. (17)). For $U/t=0.75$, there is negligible difference between the EEXX scaling method (not shown) and full EEXX DFA, which could be expected as correlation barely contributes in this regime.
Figure 2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to EEXX scaling (red) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Intermediate correlation regime – We expect this to change by turning to higher (but moderate) correlation $U/t=1.5$, results being shown on Fig. 2. For external potentials below $\Delta v=2U/t$, where the correlation energy becomes important, we see EEXX deviates from the exact result. It still improves substantially over the neglect of weight-derivative terms for $f_{-}$, but not for $f_{+}$: the weight-derivative terms only bring a small correction here, which is overestimated in EEXX. Turning to the scaling strategy, we ask: Is ground-state correlation a reasonable approximation of its ensemble analog? Results show dressing the ground-state correlation potential with a purely Hx ensemble scaling brings a modest improvement over full EEXX for $f_{-}$, but prediction of $f_{+}$ actually worsens.
These results show it is necessary to go beyond EEXX when correlation takes over. Moreover, we have to understand if the failure of Hx scaling function comes from an inherent discrepancy between ground-state and ensemble correlation, making the weight-dependent dressing strategy impractical, or if it can be fixed. To deal with situations where EEXX fails, second order (Görling-LeviGörling and Levy (1994)) perturbation theory has been applied to ensemble DFT, yielding an orbital-dependent DFA beyond EEXXYang (2021). We transpose this in the Hubbard dimer through second-order Taylor expansion in $U$:
$E^{\bm{\xi}}_{\rm c}[n]\overset{\rm PT2}{\approx}\frac{U^{2}}{2}\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}$ (18)
Compared to EEXX-only on Fig. 3, it shows much better agreement with exact Fukui functions.
Figure 3: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to PT2 (green) and neglect of weight derivative terms (grey). Exact results are shown in orange.
Moreover, comparison of functional forms of EEXX and PT2 shows where our Hx-only scaling is failing: The ratio of PT2 ensemble and ground-state correlation potentials is an explicit function of the weights and of the ensemble density:
$s_{\rm c,PT2}(\bm{\xi})=\frac{v^{\bm{\xi}}_{\rm c,PT2}[n^{\bm{\xi}}]}{v_{\rm c,PT2}[n^{\bm{\xi}}]}=s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})$ (19)
while in the Hx ratio Eq. (17) all dependence on the ensemble density simplifies out, leaving only explicit weight dependences. This disparity in complexity for the Hx and correlation part suggests using a different scaling function for each, which reads at the PT2 level:
$v^{\bm{\xi}}_{\rm Hxc}[n]\approx s_{\rm Hx}(\bm{\xi})\;v_{\rm Hx}[n]+s_{\rm c,PT2}(\bm{\xi},n^{\bm{\xi}})\;v_{\rm c}[n]$ (20)
with exact ensemble density dependence shown explicitly. Note if the correlation term was neglected all-together, leaving only the first term, the potential would be equal to that of EEXX (in the 0-weight limit). But at the PT2 level, it dresses the exact ground-state correlation with an ensemble-density dependence. The question now becomes: Does this double scaling at the PT2 level improve over the full PT2 EDFA? On Fig. 4, we compare their performance in computing $f_{+}$ for $U/t$ up to 2.5. We can see that the PT2 scaling accurately matches exact results even above $U=1.5$, when both EEXX and PT2 fail. This shows the scaling function strategy holds promise, but it is crucial to treat the Hx and correlation parts separately.
Figure 4: Fukui function $f_{+}$ in the low correlation regime for null ensemble weights (parameters in the red inset). For increasing $U$, we compare our EEXX EDFA (blue) to PT2 EDFA (green) and the PT2 scaling strategy (red). Exact results (orange) and neglect of weight-derivatives (grey) are plotted for reference.
Strong correlation – Learning from high- and low-correlation limits of the EDFT proved to be an outstanding approach for excitation energies in difficult situationsGould et al. (2023), with a versatile EDFA obtained from interpolating between the two regimes. We put it to the test for Fukui functions by building a Padé interpolation between the PT2 (low $U$) and $U/t\rightarrow\infty$ limits of the ensemble Hxc functional,
$E_{\rm Hxc}^{\bm{\xi}}[n]=a^{\bm{\xi}}[n]\,U+\frac{b^{\bm{\xi}}[n]\,U^{2}}{1+c^{\bm{\xi}}[n]\,U}$ (21)
where
$\begin{split}a^{\bm{\xi}}[n]&=\frac{\partial F^{\bm{\xi}}[n]}{\partial U}\Big|_{U=0}\\ b^{\bm{\xi}}[n]&=\frac{\partial^{2}F^{\bm{\xi}}[n]}{\partial U^{2}}\Big|_{U=0}\\ c^{\bm{\xi}}[n]&=\frac{b^{\bm{\xi}}[n]}{\gamma^{\bm{\xi}}[n]-a^{\bm{\xi}}[n]}\end{split}$ (22)
and the strong-correlation-dominant term
$\gamma^{\bm{\xi}}[n]=\lim_{U/t\rightarrow\infty}E_{\rm Hxc}^{\bm{\xi}}[n]={\rm sup}\Big\{\xi_{+},\frac{\xi_{+}-\xi_{-}}{2}+\big|n-1\big|\Big\}$ (23)
Figure 5: Fukui functions in strong correlation regime for non-zero weights (parameters in the red inset). We compare the Padé EDFA (green) to EEXX prediction (blue). Exact results (orange) and neglect of the weight-derivative terms (grey) are plotted for reference.
This EDFA performs poorly (not shown) as its discontinuous behavior (through the sup function) and density-derivative discontinuity in |n1||n-1| means it lacks cross-derivatives (in $n$ and $\xi_{\pm}$) contributions to $\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n^{\bm{\xi}}}$ that are important even in the high correlation regime. Even if Eq. (23) is formally exact in the strictly-correlated regime ($U\rightarrow\infty$), it is not representative of the strongly-correlated regime (typically $U\in[5,10]$). To remedy this, the sharp sup function and absolute value were each replaced by a smoothing function whose stiffness parameters increase with $U$, thus yielding cross derivative terms while still being compatible with the exact $U/t\rightarrow\infty$ limit. To set the stiffness parameters in a way consistent with what could be done for a real system, we only considered the symmetric case of 2-electrons Hubbard dimer, i.e. $\Delta v=0$ and $\bm{\xi}=0$, which is the analog of the uniform electron gas in this context. The stiffness parameters were set so that our Padé approximation reproduces best the (ground-state) Hxc kernel. The obtained EDFA is in principle applicable for any choice of weights, even if it was smoothed by learning from the ground-state symmetric case. It is thus most interesting to test its performance for non-zero weights.
[…]
This trend carries over to higher correlation, as shown on Fig. 6 for $U/t=10$ and $\xi=0.2$, which is remarkable as the stifness parameters smoothing the $U\rightarrow\infty$ discontinuities in $\gamma^{\bm{\xi}}[n]$ where fixed using only the ground-state, symmetric limit of the Hxc functional.
Figure 6: Fukui functions in very strong correlation regime for non-zero ensemble weights (parameters in the red inset). See Fig. 5 for labels.
Summary and outlook– After introducing the (in principle) exact expression of Fukui functions within Nc EDFT, we discussed the equivalence between weight-derivatives of the Hxc potential and derivative discontinuities in the Hxc kernel. The exact relations introduced by the Authors between individual true physical and non-interacting KS densitiesFromager (2025) and responsesDupuy and Fromager (2025) are thus a promising framework to access properties beyond excitation energies within EDFT.
By considering the conventional $N$-electron ground-state DFT limit of the theory, we showed how dressing the regular (ground-state) Hxc functional with a weight-dependent scaling function allows to recycle practical (semi-)local DFAs while accurately accounting for the ensemble Hxc potential weight-derivatives term, as illustrated on Fukui functions of the Hubbard Dimer. It was instrumental to the performance of this strategy that separate scaling functions would be applied to the Hx and correlation parts of the ground-state potential, as their ensemble analogs had starkly different dependence on ensemble weights and density. This is in line with analysis of low- and high-correlation limits of EDFTGould et al. (2023), predicting a dominant dependence on the ensemble density rather than on the weights in the latter case. When used properly, the scaling function approach extended the domain of applicability of its parent approximation by exploiting an accurate ground-state functional.
We also showed building a full ensemble DFA by interpolating between known low- and high-correlation limits performed very well, provided one paid attention to avoid discontinuous behavior in the functional. Smoothing over these discontinuities by learning from the symmetric ground state proved very effective, echoing recent work on the uniform electron gas as a basis to develop excited state density functionalsLoos (2025). These findings will foster DFA development in realistic systems. As additional strategies, one could also include exact constraints satisfied by ensembles in the design of scaling functions, such as the invariance of individual-state physical densities with respect to weights (see Eq. (8)), which is currently explored. As our formalism applies to any (neutral or charged) excitation, it can bring insights into processes that PPLB could not grasp, notably excited state reactivity through the relevant Fukui functionsAn et al. (2024). This is left for future work.
Acknowledgements - This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).